OpenStudy (anonymous):

Two small spheres with mass m = 15.7 are hung by silk threads of length L = 1.11 from a common point . When the spheres are given equal charges, so that q1=q2=q, each thread hangs at an angle = theta 23.0 degrees from the vertical. what is q?

OpenStudy (anonymous):

Since the system is in equilibrium, the electric force, gravitational force, and the tension forces must all cancel out. As a reminder, the magnitude electric force is given as follows and is the same for both balls, and in this case points outwards.\[\left|\vec F_E \right| = k\frac{q_1q_2}{r^2}=\frac{kq^2}{r^2}\]You may also have seen \(k = \dfrac{1}{4\pi\epsilon_0}\), though the forms are equivalent. Draw a well labeled free-body diagram, and decompose the forces into their \(x\) and \(y\) components to solve.

OpenStudy (anonymous):

thanks!